# SynFER: Towards Boosting Facial Expression Recognition with Synthetic Data

### v2

@@ 6 Conclusion
In this paper, we propose a synthesis data framework SynFER for facial expression recognition to address the data shortage in the field. By consolidating existing FER data and annotating with multi-modal large language models, we introduce the first facial expression-related image-text pair hybrid dataset FEText. We further propose to inject facial action unit information and external knowledge from existing FER models to ensure both fine-grained control and faithful generation of the facial expression images. To incorporate the generated images into training, we propose a diffusion-based label calibrator to help rectify the robust facial expression annotations for the synthesized images. After constructing the data synthesis methodology, we investigate the effectiveness of the synthesis data across different learning paradigms, demonstrating consistent and superior performances. We further study the scaling behavior of the synthesis data for FER.
Appendix A Appendix
A.1 Hyper-parameter Studies
Step Size $\lambda$ in Semantic Guidance. For hyper-parameter analysis, we consider five configurations of the step size $\lambda$ in semantic guidance. Due to computational resource constraints, we provide results of self-supervised learning with MoCo v3 Chen et al. (2021) on 0.2M synthetic data for pre-training and report the linear probe performances on RAF-DB Li et al. (2017). Experiment results are shown in Fig. 8. It can be seen that when $\lambda$ is relatively small, its influence on the performance is relatively small. However, as $\lambda$ continues to increase, the downstream performance is severely degraded. This is because an excessive $\lambda$ would lead to severely disrupted images, as shown in Fig. 9.
Figure 8: Hyper-parameter study on the $\lambda$ in semantic guidance. We report the linear probe performances of MoCo v3 Chen et al. (2021) pre-trained with 0.2M synthetic data on RAF-DB Li et al. (2017).
Figure 9: Visualizations of images generated with different values of $\lambda$.
A.2 Experiment Setting and Implementation Details
Self-Supervised Learning. We use the widely adopted self-supervised learning library solo-learn Da Costa et al. (2022) for experiments and follow the default settings in solo-learn for various methods. Detailed settings are shown in the tables below:
Config Pre-Training Linear Probe
SimCLR BYOL MoCo v3 SimCLR BYOL MoCo v3
batch size 64 64 64 32 32 32
optimizer Lars Lars Lars SGD SGD SGD
base learning rate 0.3 0.1 0.3 1e-3 1e-3 1e-3
weight decay 1e-4 1e-6 1e-6 1e-4 1e-4 1e-4
learning rate schedule warmup cosine warmup cosine warmup cosine step (60,80) steps (60,80) steps (60,80)
epochs 200 200 200 100 100 100
augmentation RRC RRC RRC RRC+RHF RRC+RHF RRC+RHF
Table 7: Implementation details on self-supervised pre-training. RRC and RHF denote random resize crop and random horizontal flip, respectively.
Others. As all the methods for comparisons in supervised learning (Tab. 3), few-shot learning (Tab. 5) and multi-modal fine-tuning (Tab. 5) are open-source, we thus only need to rewrite the corresponding code for dataset reading to incorporate the synthetic data. We follow the default setting in each open-source code of the compared methods.
Facial Action Unit Setting. We use pre-defined facial action unit (FAU) labels to generate images corresponding to specific facial expressions as shown below:
Facial Expression FAU
Happy AU6 + AU12
Sad AU1 + AU4 + AU15
Surprise AU1 + AU2 + AU5 + AU26
Fear AU1 + AU2 + AU4 + AU5 + AU7 + AU20 + AU26
Angry AU4 + AU5 + AU7 + AU23
Disgust AU9 + AU15 + AU16
Table 8: FAU annotations to generate specific classes of facial expression images.
Step of Performing Semantic Guidance. During the synthesis process, the total denoising steps of the diffusion model are set as 50. Semantic guidance requires backward gradient computation, which would cost a large amount of GPU hours. Thereby, we only perform semantic guidance in the latter steps, which is set as the last 5 steps of the denoising. Another reason to perform semantic guidance at the latter steps is that estimated results at early steps tend to be blurry and degraded facial images, performing semantic guidance on such images might to incorrect results.
Figure 10: Synthetic images comparison between the over-smoothing images and the natural images.
A.3 Over-smoothing of Super Resolution Training Data
As shown in Fig. 10, we provide comparisons between the over-smoothing synthetic images and the more natural ones. Due to the large amount of super-resolution data in FEText, it can be seen that solely performing fine-tuning on the entire FEText significantly degrades the realism of the images, while the proposed two-stage fine-tuning strategies in Sec. 4.2.2 could effectively prevent over-smoothing.
A.4 Limitations and future work
While the effectiveness of the proposed synthetic data framework has been demonstrated through extensive experiments, its current use is limited to augmenting the training set. A more efficient and optimized approach for leveraging synthetic data remains unexplored and warrants further investigation. Additionally, the generation process remains relatively slow, particularly when incorporating semantic guidance, which is crucial for ensuring accurate and faithful data generation. Moreover, this work focuses exclusively on facial expression recognition. However, it is important to note that the synthetic data framework has potential applications in other areas of facial affective computing, such as facial action unit detection and affective valence and arousal recognition. These avenues are left for future exploration.
A.5 FEText
More examples from FEText are shown in Fig. 11.
Figure 11: Examples from FEText.

### v3

@@ 6 Conclusion
We propose a synthetic data framework SynFER for facial expression recognition to address the data shortage in the field. We introduce the first facial expression-related image-text pair dataset FEText. We inject facial action unit and external knowledge from existing FER models to ensure fine-grained control and faithful generation of the facial expression images. To incorporate the generated images into training, we propose a diffusion-based label calibrator to help rectify the annotations for the synthesized images. After constructing the data synthesis pipeline, we show the effectiveness of synthesis data across different learning paradigms. Limitations of this work are discussed in detail in supplementary material.
7 Acknowledgments
The work was supported by the National Natural Science Foundation of China under grants no. 62276170, 82261138629, 62306061, the Science and Technology Project of Guangdong Province under grants no. 2023A1515010688, the Science and Technology Innovation Commission of Shenzhen under grant no. JCYJ20220531101412030, Open Research Fund from Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ) under Grant No. GML-KF-24-11, and Guangdong Provincial Key Laboratory under grant no. 2023B1212060076. The research reported in this publication was supported by funding from King Abdullah University of Science and Technology - Center of Excellence for Generative AI, under award number 5940.